OpenStudy (anonymous):

write the linear function that has the indicated function values f(2)=6 and f(5)=-6

OpenStudy (anonymous):

Let the linear function be: f(x) = a*x + b So: f(2) = 2a + b 2a + b = 6 ----------- 1 f(x) = a*x + b f(5) = 5a + b So: 5a + b = -6 ------- 2 Subtract both the equations: 2a - 5a = 6 + 6 -3a = 12 So a = -4 Now put it in any of the equation and find b...